OpenStudy (anonymous):

a challenge if"Y"=F(x) represents third degree polynomial function such that: F"(x)<0 when X<(-2/3) and f"(x)>0 when X>(-2/3) and the curve of the function passes by (1,6) and there exists a critical point at (-1,2) find the equation of this curve

hartnn (hartnn):

is the answer: x^3 + 2 x^2 + x + 2 ???

OpenStudy (anonymous):

@hartnm the answer is right but you have to show me the idea please

OpenStudy (anonymous):

go on hartnn :)

hartnn (hartnn):

let f(x)=a x^3 + b x^2 + cx + d since,F"(x)<0 when X<(-2/3) and f"(x)>0 when X>(-2/3), F"(x) must be 0 for X=-2/3 F"(x)=6ax+2b=0 for x=-2/3 so we get b=2a----->'W' since the critical point exist on the curve,you have two points on the curve f(x) so u get two equations like this -a+b-c+d=2------>'P' a+b+c+d=6------->'Q' now critical point is at (-1,2) implies 1st derivative of f(x) is zero at that point so 3a x^2 +2bx+c=0 at x=-1 gives 3a-2b+c=0------>'Z' you have 4 equations 4 unknowns, very easily u can find them ask if u need help in that.... @hamidic